OpenStudy (anonymous):

What is the 20th term of the following arithmetic sequence? 1/2,1,3/2,2...

Parth (parthkohli):

Common difference is \(1 \over 2\)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 0 + n({1 \over 2}) }\) Substitute n with 20, and you get the answer.

OpenStudy (anonymous):

10?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

thanks :) i need help with these: 4) Find the four arithmetic means between -21 and -36. 5) Find the four arithmetic means between 100 and 135.

Parth (parthkohli):

@clholic Just find the mean of the numbers(let's take mean as x). Then, take the mean of x and -21, also x and -36. Similar for the second one.

OpenStudy (anonymous):

A1 = -21+(-3) = -24 A2 = -24+(-3) = -27 A3 = -27+(-3) = -30 A4=-30+(-3) = -33

Parth (parthkohli):

Yes! CORRECT!

OpenStudy (anonymous):

okk Im really Stuck on this :/ Find S^25 for 1/2 + 1 + 3/2 + 2 + ...

hero (hero):

If you started with this observation: 1/2, 2/2, 3/2, 4/2.... You could have quickly realized that 20/2 or 10 was the 20th term of the sequence.

Parth (parthkohli):

...how did you even